### v1

# Scaffolding Creativity: Integrating Generative AI Tools and Real-world Experiences in Business Education

Abstract.
This study explores the integration of Generative AI tools and real-world experiences in business education. Through a case study of an innovative undergraduate course, we investigate how AI-assisted learning, combined with experiential components, impacts students’ creative processes and learning outcomes. Our findings reveal that this integrated approach accelerates knowledge acquisition, enables students to overcome traditional creative barriers, and facilitates a dynamic interplay between AI-generated insights and real-world observations. The study also highlights challenges, including the need for instructors with high AI literacy and the rapid evolution of AI tools creating a moving target for curriculum design. These insights contribute to the growing body of literature on AI in education and provide actionable recommendations for educators preparing students for the complexities of modern business environments.
copyright: noneconference: ; ; copyright: none
1. INTRODUCTION
The integration of Artificial Intelligence (AI) in education has rapidly evolved, yet its application in creative and entrepreneurial education remains underexplored. While Generative AI tools have shown promise in assisting with tasks like writing and coding (Godwin-Jones, 2022; Alharbi, 2023; Becker et al., 2023), their role in fostering creativity and business acumen in educational contexts is still is in an early stage of active research and development. While current research explores diverse applications of AI in education, including but not limited to personalized learning systems, automated assessment tools, intelligent tutoring systems(Limo et al., 2023; Fuchs, 2023; Kim et al., 2019), little has been explored in the potential to scaffold complex, interdisciplinary learning experiences.
Simultaneously, business education faces significant challenges in keeping pace with the rapidly evolving landscape of AI and its implications for the business world. This lag is exacerbated by the scarcity of pedagogical resources tailored for business students(Xu and Babaian, 2021) , contributing to what has been described as ’curriculum chaos’ (Bunch, 2020). As Sollosy and McInerney (Sollosy and McInerney, 2022) note, changing curriculum is often a ’Herculean task,’ with the rapid pace of change in the business world leading to equally rapid and sometimes disorienting changes in management education.
The interdisciplinary nature of AI also often conflicts with the traditional, discipline-by-discipline approach often employed in business schools (Currie and Knights, 2003). This creates a disconnect between the siloed nature of business education and the integrative demands of AI in business practice, potentially leaving graduates ill-prepared for the dynamic challenges of an AI-influenced business environment.
The challenge is further compounded by vast variations in students’ AI literacy (Hornberger et al., 2023) and their ability to effectively utilize rapidly evolving AI tools. Novice users, in particular, face significant hurdles in prompt designs (Zamfirescu-Pereira et al., 2023). The rapid rollout of advanced language models like GPT-4o, and Claude 3.5 Sonnet, along with image and video generation models such as Midjourney and Runway, presents a moving target for educators. While students often have some level of access to these tools (e.g., free daily messages), their proficiency in leveraging these technologies for creative and entrepreneurial tasks varies widely. Moreover, most current business educational research has focused primarily on text-based models, leaving a significant gap in understanding how to effectively integrate multi-modal AI tools into business education.
These deficiencies highlight the crucial role of instructors in guiding students’ interaction with AI. Educators must not only teach the technical aspects of AI tools but also foster the development of prerequisite skills such as critical thinking, digital literacy, and media literacy (Chiu et al., 2022). This complex educational landscape raises critical questions for instructors about the optimal level of scaffolding and how to strike the right balance between AI assistance and human creativity in the learning process. This also requires a shift towards more interdisciplinary teaching approaches (Chiu et al., 2022; Chiu, 2023) that reflect the integrative nature of AI in business.
To address these issues, this case study reports on the student learning processes and outcomes in an innovative undergraduate business course that combines AI-assisted learning with real-world experiences, through an interdisciplinary breadth-over-depth teaching approach.
We find that this integrated approach:
  1. (1)
    Accelerates knowledge acquisition and content creation, enabling students to produce high-quality outputs at an unprecedented pace.
  2. (2)
    Allows business students to overcome traditional creative skill barriers (such as design and copywriting) that previously limited their ability to produce comprehensive, multi-faceted business outputs beyond a PowerPoint Presentation.
  3. (3)
    Facilitates a dynamic interplay between AI-generated knowledge and insights and real-world observations and validation.
  4. (4)
    Complements the siloed and fragmented educational experiences.
  5. (5)
    Encourages greater expression of individual ideas in projects.
We also find further challenges:
  1. (1)
    There is now an increased demand for instructors with high AI literacy, interdisciplinary expertise, and relevant localized knowledge.
  2. (2)
    Rapidly evolving AI tools create a moving target for curriculum design, questioning the sustainability of current educational models.
  3. (3)
    Traditional pedagogical timelines and assessment methods are challenged by accelerated learning processes and breadth-over-depth approaches.
  4. (4)
    Logistical and scalability issues of immersive real-world components may limit widespread adoption.
Overall, our findings not only contribute to the growing body of literature on AI in education but also provide actionable insights for educators seeking to prepare students for the complexities of modern business environments. Furthermore, our insights challenges traditional pedagogical models, suggesting new directions for innovative interdisciplinary curriculum designs that fosters creativity and practical skills in the AI-augmented business landscape.
2. STUDY DESIGN AND METHODOLOGY
2.1. Course Overview
The course Reinventing the Brand was originally designed for undergraduate students to apply branding and user-centered design principles in reimagining existing well-known brands. However, recognizing the increasing prevalence of individual-created brands on social media and the decreasing costs of AI-assisted content creation, the course evolved to focus on students inventing their own brands. This shift also recognized the potential for AI to reshape the business landscape, potentially enabling small business owners to flourish. By having students create their own brands, the course aimed to prepare them for a future where AI might lower branding barriers to entry in various industries.
The curriculum allowed students to leverage their existing knowledge while exploring the practical challenges of brand creation in an AI-augmented environment. Students are required to produce tangible outputs including brand websites with logos and designs, social media posts, presentations detailing brand development rationales, AI-generated commercial videos, and physical artifacts such as packaging prototypes and fabric samples.
This curriculum design reflects a deliberate breadth-over-depth approach, aimed at providing students with a comprehensive overview of the multifaceted nature of brand creation and management in the digital age. The course structure is designed to break down the traditional discipline silos in business education and allow for personalized interdisciplinary learning trajectories.
2.2. Participants
The course enrolled a total of 8 students (Table 1), representing a diverse international cohort from Europe, North America, and various regions of Asia, encompassing six different nationalities.
Table 1. Participants
ID Year Major
S1 Senior Traditional Business
S2 Sophomore Interdisciplinary (Business focus)
S3 Junior Interdisciplinary (Business focus)
S4 Sophomore Interdisciplinary (Design focus)
S5 Sophomore Interdisciplinary (Design focus)
S6 Freshman Undeclared
S7 Junior Interdisciplinary (Business focus)
S8 Freshman Undeclared
  • Note: S7 and S8 completed the final project as a group.
Despite the presence of some students with business-related coursework, the overall background in business among the participants was limited. Importantly, while all students reported that they had never received any formal AI literacy training prior to this course, they exhibited varying levels of engagement experience with ChatGPT-3.5.
2.3. Phased Integration of AI Tools and Real-World Experiences
Phase 1: Text-based AI (ChatGPT-3.5)
Students were encouraged to use ChatGPT throughout the course. At the beginning of the semester, most students used ChatGPT 3.5, as GPT-4o was not freely accessible until mid-semester (capped at 5 free messages per day for the free tier). The use of ChatGPT was optional, leading to varied adoption rates among students, with many choosing not to use it at the beginning.
Table 2. Case Study Focus
Scope of Focus Details Sample size (N)
Mini-Presentation 4 (MP4) Research on raw materials and production 8
MP4 ChatGPT documentation Prompt history from three students 3
Field Trip to Yiwu Observation and reflection session 8
Branding Proposal Initial proposal for the brand concepts (text-based)a 7
Midjourney documentation Interaction history from 4 Students 4
Final Brand Portfolio Holistic analysis of the AI-generated elements 8
  • a
    Students were encouraged to submit their prompt histories but are not required, resulting in an even smaller sample size for ChatGPT and Midjourney prompt analyses.
Phase 2: Real-World Experience Integration
The course integrated two key real-world experiences to complement AI-assisted learning. Initially, students visited a local shopping mall to observe brand strategies and consumer behaviors, applying foundational business concepts to retail environments. This consumer-facing experience set the stage for understanding the broader supply chain.
The centerpiece of experiential learning was a two-day field trip to Yiwu Small Merchandise Market in Zhejiang Province, China’s hub for small commodity wholesale market, which spans 5.5 million square meters with over 75,000 shops, attracting 500,000+ international buyers yearly. This immersion exposed students to global trade dynamics and emerging consumption trends.
This experiential learning was later complemented by students re-engaging with manufacturers via online wholesale platforms. With their newly acquired knowledge and insights, students leveraged these interactions to finalize their brand development projects.
This real-world component served a dual purpose:
  • Opportunity to validate AI-generated insights: Students could test the assumptions and insights generated by AI during the local trip, creating a dynamic interplay between AI-generated output and real-world experiences.
  • Experiential learning: Students experienced firsthand the nuances of human interaction in business settings and the value of experiential knowledge – aspects that AI tools currently cannot fully replicate.
Phase 3: Multimodal AI Introduction
To facilitate the creation of tangible outputs such as brand images and AI video commercials, we introduced a Basic GenAI training session in Week 9. This session covered tools such as Midjourney, Leonardo.AI and Pika. For logo creation and voiceover, students utilized various third-party specialized AI-powered tools.
2.4. Scope of Focus and Research Methodology
This study employed a multi-phase, mixed-methods approach to investigate the integration of AI tools in business education. Despite the small sample size (N=8), the depth and richness of data collected across multiple phases allowed for a comprehensive exploration of AI-assisted learning in this context.
The course’s assignments include eight mini-presentations (MPs), a midterm branding proposal, and a brand portfolio, which includes a presentation, brand website, social media posts as well as an AI video commercial. The MPs serve the purpose of scaffolding students’ knowledge and skills to create their final culminating project.
Given the exploratory nature of the study and the diverse data collected, our analysis focused on three key areas (Table 2):
  1. (1)
    AI-assisted learning in unfamiliar domains: We examined MP4 outputs and available prompt histories ($N=3$) to understand how students used ChatGPT to research unfamiliar topics.
  2. (2)
    Integration of AI insights with real-world experiences: We compared students’ AI-generated knowledge before the Yiwu trip with their post-trip understanding. This analysis was based on observational notes from the field trip and discussions with students during and after the visit.
  3. (3)
    AI-assisted visual transformation: We analyzed the evolution of students’ brand concepts from text-based proposals to visual portfolios. This included examining Midjourney interaction history ($n=4$) and comparing initial brand proposals with final visual presentations.
Limitations of Small Sample Size
Although students were highly encouraged to submit their entire prompt histories, we received only a handful of submissions. While this small sample size limits our ability to draw broad conclusions from the prompt analysis, this study serves as a foundation for future, larger-scale investigations into the integration of AI tools in educational settings.
3. FINDINGS
Our analysis focuses on three components of the course: Mini-Presentation 4 (MP4), the Yiwu field trip, and students’ use of Midjourney for their final presentations and AI video commercials. These components provide many insights into how students leverage text-based AI tools for research, integrate real-world experiences, and use multimodal tools as creative scaffolds for visual transformation.
Table 3. Student Prompts and Observed Behaviors
Prompts Observed Behavior
S1 begins with specific materials
P1 Which items of clothing can be made with 100% hemp fabric? And with 100% Tencel? Demonstrates initial domain knowledge
P2 What differences are there in the production of cotton jeans and hemp jeans from a sustainability perspective? Narrows down; shows curiosity
P3 What are the most eco-friendly fibers for clothing? Returns to broader perspective
P4 List the advantages of organic cotton, hemp and tencel Focuses on specific materials
P5 Are all of these fabrics biodegradable? Confirmatory question
P6 What are the basic production processes of creating hemp, cotton and tencel fabrics? And what are the basic steps to turn them into clothing? Shifts focus to production processes
P7 Explain how ramie is a sustainable fabric, as well as its advantages and disadvantages Introduces new material (tangential)
P8 How do you produce ramie fabric? Follows up on new material
S2 begins with design-related prompts
P1 I want to make a tote bag designed for students, but can also be used for casual outings, with water bottle pockets and … Demonstrates initial domain knowledge; design goals
P2 What materials are commonly used for tote bags Returns to broader perspective
P3 Which material choice would be most durable, water proof and more suitable for student Compare options
P4 The internet shows that nylon and polyster blend fabrics are recommended; what do you think External source validation
P5 What material can be used for the water bottle pockets, since it has to be flexible and stretchy for various sizes of water bottles Focus on specifics
P6 What are the basic production steps for a bag Shifts focus to production process
S3 adopts a more broad to specific approach
P1 What materials are mainly used for black/textured hoodies? Starts with broad question
P2 What is the raw material used for graphic designs on hoodies? Shifts focus to specific component
P3 How are plasticol ink and heat transfer vinyl used together? Shifts focus to specific materials
P4 Can you list the steps of manufacturing hoodies with graphic designs? Shifts focus to production process
  • 1
    Students were asked to identify the raw materials and the production process of their brand merchandise to ensure that they are aligned with their brand designs. Hence all students shifted to production process at one point in their prompt histories.
  • 2
    Students with prior domain knowledge (S1 and S2) exhibited non-linear learning patterns, exploring both specific and broad concepts iteratively. All three students showed accelerated acquisition of domain-specific knowledge.
3.1. AI-Assisted Research in Unfamiliar Topics
In MP4, students investigated production processes for their brand concepts to understand how product features are affected by raw materials, production and supply chain dynamics, which are not typically taught in business programs. This assignment was strategically designed to observe how students utilize ChatGPT to research unfamiliar topics typically not covered in traditional business curricula.
The assignment instructions included general ChatGPT usage principles:
  • encouraged open-ended exploration
  • start with broad questions to establish a foundation
  • take an iterative approach
An additional aspect of the assignment instructions was the deliberate de-emphasis on accuracy. Unlike typical academic requirements, students were explicitly told that “100% accuracy” was not the goal. This approach aimed to encourage engagement with unfamiliar topics and facilitate broader exploration without the constraints of constant fact-checking when time is limited.
All of the students used ChatGPT-3.5 to aid them in their research process, but only 3 out of the 8 students documented their entire prompt histories in their submissions, all of which exhibited adaptive prompting skills.
Analysis of the documented prompt histories and final presentations revealed several interesting patterns:
Non-Linear Learning Trajectories
Despite the assignment instructions being structured in a typical linear format – from broad to narrow – to help students narrow down their research focus, the sequencing of the questions some students asked revealed non-linear learning trajectories. This divergence from the expected pattern suggests that AI-assisted research might fundamentally alter how students navigate unfamiliar topics.
Students with strong domain knowledge in certain fields demonstrated distinct starting points in their AI interactions. For instance, S1, leveraging prior internship experience in garment production, began with questions of the raw materials she already knew were eco-friendly, by asking what garments can be made from these materials. This approach indicates an attempt to validate and expand existing knowledge using ChatGPT as a reference point.
In contrast, S2 drawing on design experience, initiated the interaction by describing her product vision with clear design-oriented specifications such as “water bottle pockets” and “enough space for students”. Possibly due to the lack of experience with ChatGPT, S2 did not ask any question after specifying her product requirement. Realizing that ChatGPT’s response was not what she was looking for, she returned to a more general question of what materials are usually used for her product categories.
S3, who had limited experience in either production or design, followed a more general and linear questioning pattern in the prompt history. This pattern suggests that students without prior domain knowledge may rely more heavily on the structured approach provided in the instructions.
Students’ background in domain knowledge also seemingly affected their analytical approach when interacting with ChatGPT. Students with less domain knowledge tend to accept AI responses more readily, using them as scaffolding to build general understanding. But the students with certain domain knowledge are more likely to demonstrated a more critical and nuanced engagement with AI-generated information.
For example, when S1 learned that hemp can be used to create “even jeans” she immediately followed up with a question about the differences between cotton and hemp jeans for deeper inquiry.
Similarly, S2 cross-referenced AI responses with internet sources, particularly when ChatGPT’s answers seemed too theoretical for her practical, production-oriented questions. This triangulation between AI, web resources, and personal knowledge illustrates a sophisticated approach to information validation that leverages domain expertise.
the internet shows that nylon and polyster blend fabrics are recommended, what do you think? [S2]
Moreover, one of the students demonstrated a tendency to inquire about tangential knowledge. In final queries, S1 focused on ramie, a fabric not mentioned in previous responses. This sudden shift in focus could represent either a recollection of prior knowledge or genuine curiosity sparked by the AI interaction. This behavior could suggest the AI’s responses might trigger memories of related concepts or introduce new ideas that pique the student’s interest, leading to unexpected learning paths.
Dynamic Shift of Scope of Questioning
Throughout their interactions with ChatGPT, students demonstrated a notable pattern of oscillating between broad and narrow inquiries. This dynamic shift in the scope of questioning reveals a unique aspect of AI-assisted learning.
For instance, S1 began with a specific question about eco-friendly fabrics, then broadened her inquiry to ask about ”the most eco-friendly fibers for clothing,” before narrowing again to inquire about specific properties of organic cotton, hemp, and Tencel.
Similarly, S2 started with a broad product description, narrowed to specific material inquiries, then shifted the focus to general production processes, which is the second part of the assignment.
S3 added her own comments in her prompt history documentation, mentioned she “had no idea there were a variety of ways to produce hoodie’s graphic designs” after P2. ChatGPT’s response, presented as a long list of bullet points, initially led her to incorrectly assume that some of the materials are used the same processes. Upon realizing her misunderstanding from follow-up responses, she broadened her inquiry to ask about the general production process of hoodies.
This oscillation between broad and narrow questioning reflects an adaptive learning which allows the students to gain contextual understanding and specification of details. There is also seemingly redirection. Students also tend to redirect their focus or scope after getting responses with unexpected information or concepts.
These cases may suggest that AI-assisted learning could allow students to rapidly shift between scopes to facilitate a more flexible and responsive research process. However, this pattern also presents challenges. The ease of shifting scope might lead to a scattered learning process if not managed well.
Accelerated Domain-Specific Knowledge Acquisition
Another interesting observation across all eight students’ presentations was the rapid development of project-specific knowledge in product development and manufacturing. The prompt histories from the 3 samples revealed that students could navigate technically-oriented questions and answers and complete their research process within 4-8 prompts.
This observation opens up exciting possibilities for personalized, project-based learning, where students can quickly acquire the specific expertise needed for their individual projects within compressed timelines. It also highlights the potential of AI as a powerful tool for just-in-time learning in professional contexts.
While it’s understandable that business students researching unfamiliar disciplines aren’t necessarily required to understand all the details, this rapid acquisition of knowledge still raises important considerations:
1. What role does critical evaluation play in this accelerated process?
2. How does this rapid learning affect knowledge retention?
3. How can we contextualize the tradeoff between the speed of knowledge acquisition and the development of deeper understanding?
These questions highlight the need for further research into the cognitive processes involved in AI-assisted learning and the development of pedagogical strategies.
3.2. Dynamic Interaction with Real-World Experiences
To investigate how AI can serve as more than just an information source and understand how human-AI collaboration works in learning contexts, we designed the course specifically to bridge the gap between the AI insights and practical application in ways previously unexplored in traditional business education.
Expanding AI Capabilities in Experiential Learning
After their AI-assisted research into raw materials and production processes, students successfully identified potential manufacturers on 1688.com, a major B2B e-commerce platform. This demonstrates how AI can expand human capabilities in real-world business scenarios, allowing novices to engage in complex market research tasks typically reserved for experienced professionals.
Uncovering AI Limitations through Real-World Engagement
To further uncover the practicality of the AI-generated knowledge and synethsize it into the entire experiential learning, the students went to Yiwu Small Merchandise Market to gain firsthand market intelligence, and have a chance to have a real-world validation in AI-assisted learning.
Not surprisingly, not everything what ChatGPT provided proved to be useful in reality. For example, S1’s discovery that 100% hemp production was far less common than AI had implied. Moreover, many students encountered a wide array of materials and production techniques during their visit that were not mentioned in their AI-assisted research.
This discovery points to two known limitations of AI-generated knowledge: first, its tendency to provide generic answers unless prompted with very specific questions, which is difficult for student users; and second, its potential lag in incorporating the latest innovations or highly specialized industry knowledge.
It suggests a need for AI interfaces that encourages a continuous learning loop, where real-world experiences and additional knowledge sources not only supplemented AI-generated information but also informed future AI interactions.
Human-AI Synergy: Navigating Complex Business Environments
Perhaps the most significant insight emerged from the unexpected ways students combined AI-generated knowledge with real-world experiences to navigate the opaque Yiwu market. Contrary to expectations, the rapid knowledge acquisition facilitated by AI allowed some students to successfully and confidently pose as experienced buyers during vendor interactions, gaining valuable market intelligence.
This phenomenon was particularly notable given that the vendors can easily identify novice business buyers or ”tourists” (retail buyers) if they don’t ask the ”right” questions, and offer unfavorable deals. The AI-assisted preparation appeared to provide students with sufficient domain knowledge to ask pertinent questions and engage in more nuanced conversations, potentially altering vendor perceptions and interactions.
The observed synergy between AI preparation and human adaptability suggests a novel model of human-AI collaboration in learning contexts. Students utilized AI-generated knowledge as a foundational framework, which they then dynamically adjusted based on real-time observations and interactions in experiential learning. This process of continuous recalibration between AI inputs and human experiences suggests a more fluid and interactive model of AI-assisted learning than previously conceptualized.
The opacity of the Yiwu market, with its unwritten rules and subtle negotiation cues, highlighted areas where human experiential learning still surpasses AI capabilities. However, the AI-assisted preparation enabled students to engage more confidently with these tacit aspects of business interactions, suggesting that AI can play a crucial role in preparing learners for complex, real-world scenarios.
3.3. AI-Assisted Visual Transformation
Traditionally, business school assignments have predominantly relied on text-based presentations supplemented with basic visuals. This approach, partly attributed to students’ non-design backgrounds, has limited the exploration of visual communication in business education. Our case study investigates how the introduction of AI-assisted text-to-image and image-to-video tools impacts students’ ability to visualize and communicate business concepts.
Four students (N=4) kept record of their interaction with Midjourney in the class Discord channel after the training session. Other participants utilized various AI image generation tools through personal accounts, such as Leonardo.ai, which precluded consistent data collection across the entire cohort. Table 4 presents the statistics of these interactions, including initial concept prompts, iterative prompts (usually style or component changes), upscaling actions, remixes, uploads, image-to-text descriptions and command error corrections.
Table 4. Midjourney Interactions (Post-Training)
ID S2 S3 S4 S5
Interactions 90 31 40 12
Concepts 14 10 12 4
Iterations 45 10 19 6
Effective Interactions per Concept 2.6 1.0 1.6 1.5
  • 1
    Students could make errors in their prompts due to their unfamiliarity with the command format in Midjourney’s server on Discord, causing ineffective interactions.
AI as a Cognitive Scaffold and Human-AI Co-Creation
Our study revealed that AI tools functioned not merely as implementation aids, but as collaborative partners in both concept development and visualization. This partnership manifested as an iterative process where AI-generated visuals influenced conceptual thinking, which in turn guided further visual exploration.
In particular, the introduction of multimodal AI tools after midterm appeared to serve as cognitive scaffolds, enabling students to bridge the gap between abstract business concepts and concrete visual representations. This scaffolding effect was evident in the expansion and refinement of initial ideas:
Concept Expansion and Refinement: Visual presentations often expanded upon initial text-based ideas, suggesting that the AI-assisted visual creation process prompted students to further refine and develop their brand concepts.
Nuanced Expression: The visual medium allowed students to express nuanced ideas that were difficult to convey through text alone, indicating that AI tools may be facilitating more complex conceptual thinking.
For example, S3’s brand concept aimed to ”cross the line between femininity and masculinity,” a nuanced idea difficult to convey through text alone. Figure 1 showcases the concept images S3 generated in an iterative cycle on Midjourney for her brand portfolio and AI commercial. In post-class correspondence, S3 detailed her iterative prompting process:
”… was trying to display feminine hands doing things that typically men are associated with the most: such as playing the drums and fixing a car”. But because PikaLabs she used had a tendency to generate extra fingers, she “turned to generating images of feminine hands fixing a car.”
This iterative process led S3 to have deeper reflections on her brand’s message and target audience:
”I realized my initial concept might appeal to an unintended audience…I generated images of a man and a woman fixing cars together to try to promote what I believe is a better envisionment of gender harmony/equality”
Figure 1. S3’s Interactions with Midjourney Reflect Brand Concept Refinement
1st concept: a first person POV realistic picture of a pair of hands with long pink nails playing the drums
2nd concept: a first person POV realistic picture of a pair of hands with long pink nails holding a screwdriver while fixing a car’s engine
3rd concept: a realistic photo shoot of a woman with orange hair holding a screwdriver while happily helping her guy friend fix his car
The iterative visual exploration not only helped S3 refine her brand concept but also prompted a more nuanced refinement of her target audience definition and brand values.
Moreover, analysis of prompt histories revealed that students often incorporated unexpected elements from AI generations into their subsequent prompts. Initially focusing on product features in her prompts, S2 noticed a blurry figure in one AI-generated image. This unexpected element inspired her to incorporate human figures in subsequent prompts, which are better representation of her easy-to-use brand concepts.
Analysis of S4’s prompt history revealed significant variations in length and complexity ($\mu=51.23$, $\sigma=30.74$). This variability suggests potential utilization of text-based AI tools or prompt aid tools to generate Midjourney prompts. This observation indicates that even novice users are capable of leveraging multiple AI tools in tandem to enhance their creative process.
This collaborative process aligns with the vision that AI tools act not just as implementers of human ideas, but as active collaborators in the creative process. It also resonates with the notion of AI as a coach or a colleague (Lubart, 2005), stimulating new directions of thought and encouraging students to reconsider and enhance their initial concepts.
Democratization of Visual Design Skills
The rapid adaptation to visual communication also suggests that AI tools may be democratizing visual design skills in non-design disciplines. All students demonstrated a strong capability to shift towards visual communication, regardless of their prior design experience.
This democratization was particularly evident in the professional-quality outputs produced by the students. The quality and sophistication of these visual materials markedly exceeded typical outputs observed in traditional business education courses, particularly among first- and second-year students.
As noted previously, none of the students had prior experience with text-to-image or image-to-video tools in an academic or professional context before the in-class training session in Week 9. While some students may have casually experimented with these tools for personal amusement, they had not applied them to serious project work before this course. This rapid skill acquisition in a professional context suggests that students may possess an innate capacity for visual expression that was previously untapped in traditional business education. AI tools appear to be unlocking this potential, raising important questions about the future of visual literacy in business education.
Shift Towards Individual Expression and Engagement
The introduction of AI tools in the course also led to an unexpected shift in student behavior, challenging traditional collaborative norms in business education. Despite encouragement to form pairs after the brand proposal stage, 6 out of 8 students insisted on working individually on their projects. This preference for individual work, especially given the comprehensive nature of creating an entire brand portfolio, marks a significant departure from typical business education practices.
This shift towards individual work aligns with the concept of ”AI as an Empowering Partner” (UNESCO, 2023). The AI tools appeared to increase students’ sense of agency and self-efficacy, enabling them to tackle complex design tasks that would typically require team effort or specialized skills.
The high motivation observed in students to express and implement their ideas independently may also resonates with Csikszentmihalyi’s Flow Theory(Csikszentmihalyi and LeFevre, 1989). The AI tools seem to have created an optimal challenge-skill balance, facilitating a state of flow where students felt both capable and motivated to pursue ambitious individual projects.
Moreover, while traditional business education often relies on human-human collaboration to compensate for individual skill gaps, our study suggests that AI tools are redefining this dynamic. This phenomenon can be understood through the lens of Human-AI Collaboration [16]. This also paves the way for a new pedagogical shift in personalized learning.
4. IMPLICATIONS AND LIMITATIONS
4.1. Implications for Business Education
Our case study reveals a fundamental shift in the landscape of business education, one that demands a radical rethinking of current pedagogical models. The integration of AI tools, coupled with real-world experiences, points towards a future where business education transcends traditional disciplinary boundaries, becoming inherently interdisciplinary and deeply intertwined with technological literacy.
The observed non-linear learning trajectories and rapid skill acquisition facilitated by AI tools suggest that the future of business education lies not in the traditional transmission of static knowledge, but in the development of adaptive learning capabilities. This shift addresses the ’curriculum chaos’ noted by [10] by embracing flexibility as a core principle. Rather than struggling to keep pace with rapidly changing business landscapes, educational institutions should focus on creating dynamic learning environments where AI serves as a collaborative partner in the educational process. This approach could potentially resolve the tension between the need for current, relevant knowledge and the practical constraints of curriculum development.
Furthermore, the democratization of design and technical skills through AI tools points to a future where the boundaries between business, design, and technology become increasingly blurred. This convergence calls for a reimagining of business education as a holistic, integrative discipline. Future business curricula should not merely incorporate AI as a tool, but should be fundamentally restructured around the principle of human-AI collaboration. This could involve the development of new hybrid courses that blend elements of business strategy, design thinking, and AI interaction, preparing students for a world where these elements are inseparable in practice.
The shift towards individual expression enabled by AI tools, coupled with the continued importance of real-world collaborative experiences, suggests a need for new models of collaborative learning. Future business education should explore AI-mediated collaboration platforms that can balance individual creativity with team-based problem-solving. This could potentially address the challenge of developing collaborative skills in an AI-augmented world, where individual capabilities are significantly enhanced by technology.
Perhaps most critically, our findings underscore the need for a new approach to developing business acumen in the AI era. Rather than focusing on the acquisition of specific skills or knowledge sets, which may quickly become obsolete, business education should prioritize the development of ’AI-interaction literacy’. This involves not just technical proficiency with AI tools, but a deep understanding of how to critically evaluate AI-generated insights, how to effectively prompt and guide AI systems, and how to synthesize AI-produced information with human expertise and real-world observations.
The integration of real-world experiences with AI-assisted learning in our study also points to a future where the boundaries between classroom and field blur. Business schools should consider adopting more fluid educational models where AI-enhanced project-based learning is seamlessly integrated with real-world business experiences. This could involve AI-guided virtual internships, augmented reality business simulations, or AI-facilitated global collaboration projects.
4.2. Limitations and Challenges
First, the small sample size and single course context limit the generalizability of our results. Future studies should expand to larger cohorts across multiple courses and institutions to validate and extend our findings. This expansion would also help address potential self-selection bias, as students who chose to take this innovative course might be more receptive to AI integration than the general student population.
The lack of longitudinal data is another limitation. While we observed rapid skill acquisition and knowledge development, the long-term retention and application of these skills remain unexplored.
The potential novelty effect of AI tools is also a consideration. The heightened engagement and outcomes we observed might partially result from the novelty of the technology. Longitudinal studies are needed to determine if these positive effects persist as AI tools become more commonplace in educational settings.
Assessment challenges in an AI-assisted learning environment emerged as a significant limitation. Distinguishing between AI-generated content and student-generated insights proved difficult, raising questions about how to fairly evaluate learning outcomes. This challenge calls for the development of new assessment methodologies tailored to AI-integrated curricula, an important area for future research.
The scalability of our approach presents challenges beyond mere technical implementation. A striking example emerged when a student, aiming to generate images representing eco-friendly garment production, consistently received AI-generated visuals of Asian factory settings with Asian workers. This outcome, while reflective of current global manufacturing trends, inadvertently highlighted the intricate interplay between AI biases, global economic realities, and brand representation.
This highlights the necessity to develop critical thinking skills that transcend traditional business silos, requiring both educators and students to analyze AI-generated insights within broader social, ethical, and economic contexts.
In conclusion, the future of business education should be dynamic, AI-enhanced learning ecosystems that prioritize adaptability, interdisciplinary integration, and the development of new forms of literacy. This vision challenges institutions to move beyond incremental changes and towards a fundamental reimagining of what business education can and should be in an AI-augmented world. For the HCI community, this opens up new avenues of research into the design of AI systems that can support this educational transformation, from adaptive learning interfaces to AI-mediated collaboration tools.